OpenStudy (anonymous):

find the slope intercept for the line satisfying the following conditions x intercept 2, y intercept 2/3

OpenStudy (anonymous):

there is something missing .. unless it is assumed that the line passes through 0,0 as well or something

OpenStudy (anonymous):

thats all thats in the question

OpenStudy (ghazi):

well consider point (2,0) and (0,2/3) can you write it now?

OpenStudy (ghazi):

y-0=(2/3-0)/(0-2) * (x-2)

OpenStudy (anonymous):

well i guess ghazi right about it ..

OpenStudy (ghazi):

solve it you'll have equation of line

OpenStudy (ghazi):

i guess that would be y=-1/3 *(x-2)

OpenStudy (anonymous):

so thats the slope intercept form

OpenStudy (ghazi):

yes it is y=-1/3 x+2/3 your slope = -1/3

OpenStudy (anonymous):

thank you both

OpenStudy (ghazi):

:) YW